A capacitor is constructed from two conductive metal plates 30cm x 50cm


which are spaced 6mm apart from each other, and uses dry air as its only


dielectric material. Calculate the capacitance of the capacitor. (0.221nF).

1
Expert's answer
2022-03-15T10:33:48-0400

The capacitance of the parallel plates capacitor

"C=\frac{\epsilon_0A}{d}"

"C=\frac{8.85*10^{-12}*0.3*0.5}{0.006}=2.21*10^{-10}\:\rm F=0.221\: n F"